OpenStudy (anonymous):

derivative of 6sin^2x note: only the sin is ^2. the x is not.

OpenStudy (anonymous):

42

OpenStudy (anonymous):

lol...

hartnn (hartnn):

whats derivative of x^2 ?

OpenStudy (anonymous):

2x!

OpenStudy (anonymous):

its 42

hartnn (hartnn):

so what will be derivative if sin^2 x ??

OpenStudy (anonymous):

12sinx? ._.

hartnn (hartnn):

now use chain rule d/dx 6sin^2 x = 12 sin x d/dx sin x so whats derivative of sin x ?

OpenStudy (anonymous):

cosx

hartnn (hartnn):

yes :) so d/dx 6sin^2 x = 12 sin x . cos x can u simplify this further ?

OpenStudy (anonymous):

No, dont think so.

hartnn (hartnn):

u know formula for sin 2x ?

OpenStudy (anonymous):

nope

hartnn (hartnn):

sin 2x = 2 sin x cos x so 12 sin x cos x = 6 *(2sin x cos x) = 6 sin 2x. d/dx 6sin^2 x = 12 sin x . cos x = 6 sin 2x got this ?

OpenStudy (anonymous):

i think the answer 12 sin x . cos x is good for my purposes. dont need to simplify any more than that i think. ty

hartnn (hartnn):

ok. welcome :)